# QueST: Self-Supervised Skill Abstractions for Learning Continuous Control

### v2

@@ 3.1 Problem Setting
We consider a dataset where is a continuous-valued action and is a tuple consisting of a high-dimensional sensory observation. The data is collected via either human teleoperation or scripted policies for M different task each with a label . In our setting, consist of RGB image observations from the front camera and gripper camera (if available) along with proprioceptive state of the agent. is a natural language description of the task but can also be a one-hot encoding.
@@ 3.2 Finite Scalar Quantization
Given a feature vector to quantize, instead of learning a parameterized codebook [64] and quantizing by matching the nearest neighbor in the codebook, FSQ quantizes by first bounding it into certain range with (e.g. ) and then rounding each dimension into integer numbers directly. defines the width of the codebook for each dimension (e.g. ). Finally, it is easy to see that the size of the quantization space is . An MLP can be used to transform into continuous space of required dimension further.
@@ 5 Experiments
We design the experiments to empirically evaluate the performance of Quantized Skill Transformer in three practical settings: (1) Multitask IL, (2) Few-shot adaptation, and (3) Long-horizon IL. Lastly, we perform some ablations to empirically justify our model design choices.
@@ 5.1 Benchmarks and Baselines
LIBERO [37] is a lifelong learning benchmark featuring several task suites consisting of a variety of language-labeled rigid- and articulated-body manipulation tasks. Specifically, we evaluate on the LIBERO-90 suite, which consists of 90 manipulation tasks, and the LIBERO-LONG suite, which consists of 10 long-horizon tasks composed of two tasks from the LIBERO-90 suite. As described in more detail below, we use the LIBERO benchmark to study the multitask IL, few-shot adaptation and long-horizon IL settings. Because tasks from this benchmark are language-annotated, we use the output of a frozen CLIP [53] encoder for the task conditioning input .
@@ 5.2 Performance on Multitask BC
We evaluate the goal-conditioned multi-task imitation learning capabilities of QueST and the baselines using the LIBERO-90 and ML45 benchmark suites. For LIBERO-90, the learner receives 50 expert demonstrations per task from the author-provided dataset. For ML45, we use the provided scripted policies to collect 100 demonstrations per task. We evaluate the model at the end of training and for each task run 40 evaluation rollouts (50 for MetaWorld) starting from the initial states selected sequentially from a predefined set. We report the aggregated results across 4 seeds (5 seeds for MetaWorld).
In Figure 2(a), we present the average success rate across 90 tasks in LIBERO-90 against the aforementioned baselines. Quantized Skill Transformer achieves state-of-the-art results on LIBERO-90 benchmark, outperforming the best-performing baseline (VQ-BeT) by a margin of 8%. We attribute its performance to its learned latent space that enables effective knowledge sharing across tasks. While VQ-BeT also shows strong performance, we see that QueST’s architecture lends itself better to sharing representations across tasks. Overall all LVM baselines outperform ResNet-T, a simple BC baseline, suggesting that it’s a more effective modeling approach especially in a low-data regime like robotics. Figure 3(a) shows the average success rate across 45 tasks in ML45 benchmark. Being a simpler benchmark, all methods perform almost similar in Multitask-IL setting which is consistent with the trend observed in [73].
@@
5.3 Few-shot Adaptation to Unseen Tasks
@@
Figure. 2(b) shows the average success rate for 5-shot IL across 8 unseen tasks in LIBERO-LONG. QueST achieves SOTA performance, surpassing all other baselines by an absolute margin of %. Though we see a drop of % without decoder finetuning, it still outperforms all the baselines. These results highlight the superiority of QueST in learning transferable representations of action abstractions and effectively leveraging them for downstream decision making. For a fair comparison, we also tried fine-tuning the decoder of VQ-BeT but did not observe any gains from it. VQ-BeT struggles in this setting as it heavily relies on offset head to output continuous action corrections which requires more data-samples for sufficient coverage in the continuous action space. Figure 3(b) shows the average success rate for 5-shot IL across 5 unseen tasks in MetaWorld. Similar to multitask results, all methods perform comparably, with QueST showing a slight improvement over the others. QueST leverages its learned skill tokens to compositionally model their distribution for an unseen task in just 5 demonstration examples. For few-shot evaluation protocol please refer Appendix E.1.
@@ 5.5 Ablations
Vector Quantization: We replace the FSQ layer with a Vector Quantization layer of nearly the same codebook size.
@@
Mirrored Decoder: Following a typical autoencoder design, we use a decoder that mirrors the encoder, using transposed convolutions instead of strided convolutions, and with the strides in reverse order as in the encoder. This decoder directly takes skill-token embeddings as input and outputs the continuous actions.
@@
Causality: We ablate the use of causal layers in various parts of our network.
We report the ablation results in both multitask and few-shot settings on LIBERO in Table 1 and Table 2. We see that QueST outperforms all ablations, validating our design choices.
We also perform a sensitivity experiment over several hyperparameters including downsampling factor and codebook size in Figure 4. Across the board we see that the hyperparameters are more important in the difficult few-shot learning setting. In Figure 4(a) we see that both algorithms have the best performance with a modest downsampling factor of , and in Figure 4(b) we see that QueST does well with a 1024 codebook vectors. For more discussion on ablations please refer Appendix C
@@ 6 Conclusion
This paper works towards the broader goal of automating a wide range of manipulation tasks. While this can have positive impacts, such as helping people with mobility impairments or performing dirty tasks humans would rather not do, it can also have negative impacts such as automating peoples’ jobs away and further concentrating wealth in the hands of a handful of companies. It is important that we in the machine learning community advocate for equitable use of the technology we develop.
@@ B.3 Baseline Implementation:
To ensure fair comparison of different model architectures, we use same input modalities and same observation & task encoders for all baselines. VQ-BeT needs a goal image, we instead give it task embedding as goal. Same as QueST, we concatenate observation embeddings for all modalities at any timestep and project them to respective model’s hidden dimension.
@@ Appendix C Discussion on Ablations
Appendix D Skill-space visualization
We present a t-SNE visualization (Figure 5) illustrating the learned skill-space across multiple set of similar tasks. We consider four different combinations of similar tasks to effectively examine the z-embeddings corresponding to their trajectories. Each data point in the plot represents a vector of z-embeddings at a specific timestep throughout the entire episode, with decreasing transparency indicating temporal progression. We show that the QueST encoder learns a semantically meaningful skill-space that encodes shared representations of similar motion primitives across different tasks. Notably, the skill-space learning happens in the first stage training which does not make use of any task labels.
Appendix E Additional Results
E.1 Fewshot IL
@@
In finetuning phase, we finetune ResNet-T, VQ-BeT & QueST for 100 epochs and ACT & Diffusion Policy for 200 epochs. For each task in MetaWorld, we evaluate each method across 10 evenly spaced checkpoints for 5 seeds on 50 distinct initial states and report the results corresponding to the best performing checkpoint. For Libero, we found the final checkpoint to perform best for all methods and hence report results corresponding to it across 4 seeds.

### v3

@@ 3.1 Problem Setting
We consider a dataset consisting of robot interaction trajectories where is a continuous-valued action and is a tuple consisting of a high-dimensional sensory observation. The data is collected via either human teleoperation or scripted policies for M different task each with a label . In our setting, consist of RGB image observations from the front camera and gripper camera (if available) along with proprioceptive state of the agent. is a natural language description of the task but can also be a one-hot encoding.
@@ 3.2 Finite Scalar Quantization
Given a feature vector to quantize, instead of learning a parameterized codebook [65] and quantizing by matching the nearest neighbor in the codebook, FSQ quantizes by first bounding it into certain range with (e.g. ) and then rounding each dimension into integer numbers directly with straight-through gradients (round_ste). defines the width of the codebook for each dimension (e.g. ). Finally, it is easy to see that the size of the quantization space is . An MLP can be used to transform into continuous space of required dimension further.
@@ 5 Experiments
We design the experiments to empirically evaluate the performance of Quantized Skill Transformer in three practical settings: (1) Multitask IL, (2) Few-shot transfer, and (3) Long-horizon IL. Lastly, we perform some ablations to empirically justify our model design choices.
@@ 5.1 Benchmarks and Baselines
LIBERO [38] is a lifelong learning benchmark featuring several task suites consisting of a variety of language-labeled rigid- and articulated-body manipulation tasks. Specifically, we evaluate on the LIBERO-90 suite, which consists of 90 manipulation tasks, and the LIBERO-LONG suite, which consists of 10 long-horizon tasks composed of two tasks from the LIBERO-90 suite. As described in more detail below, we use the LIBERO benchmark to study the multitask IL, few-shot transfer and long-horizon IL settings. Because tasks from this benchmark are language-annotated, we use the output of a frozen CLIP [54] encoder for the task conditioning input .
@@
For a detailed discussion between QueST and the baseline methods, please refer to Appendix C.
@@ 5.2 Performance on Multitask BC
We evaluate the goal-conditioned multi-task imitation learning capabilities of QueST and the baselines using the LIBERO-90 and ML45 benchmark suites. For LIBERO-90, the learner receives 50 expert demonstrations per task from the author-provided dataset. For ML45, we use the scripted policies provided in the official Metaworld codebase to collect 100 demonstrations per task. We evaluate the model at the end of training and for each task run 40 evaluation rollouts (50 for MetaWorld) starting from the initial states selected sequentially from a predefined set. We report the aggregated results across 4 seeds (5 seeds for MetaWorld).
In Figure 2(a), we present the average success rate across 90 tasks in LIBERO-90 against the aforementioned baselines. Quantized Skill Transformer achieves state-of-the-art results on LIBERO-90 benchmark, outperforming the baseline VQ-BeT and Diffusion Policy by a margin of 8 and 13% respectively. We attribute its performance to its learned latent space that enables effective knowledge sharing across tasks. While VQ-BeT also shows strong performance, we see that QueST’s architecture lends itself better to sharing representations across tasks. Our implementation of ResNet-T achieves significantly better performance than the reported number (16.8%) in [38] but is still lower than QueST. Figure 3(a) shows the average success rate across 45 tasks in ML45 benchmark. Being a simpler benchmark, all methods perform almost similar in Multitask-IL setting which is consistent with the trend observed in [74].
@@
5.3 Few-shot Transfer to Unseen Tasks
@@
Figure 2(b) shows the average success rate for 5-shot IL across 8 unseen tasks in LIBERO-LONG. QueST achieves SOTA performance, surpassing all other baselines by an absolute margin of %. Though we see a marginal drop of % without decoder finetuning, it still outperforms all the baselines. These results highlight the superiority of QueST in learning transferable representations of action abstractions and effectively leveraging them for downstream decision making. For a fair comparison, we also tried fine-tuning the decoder of VQ-BeT but did not observe any gains from it. VQ-BeT struggles in this setting as it heavily relies on offset head to output continuous action corrections which requires more data-samples for sufficient coverage in the continuous action space. Figure 3(b) shows the average success rate for 5-shot IL across 5 unseen tasks in MetaWorld. Similar to multitask results, all methods perform comparably, with QueST showing a slight improvement over the others. QueST leverages its learned skill tokens to compositionally model their distribution for an unseen task in just 5 demonstration examples. For few-shot evaluation protocol please refer Appendix D.1.
@@ 5.4 Long-horizon BC
5.5 Latency
Our pipeline runs at 33Hz, which is more than suffcient for vision-based real-robot control where most camera systems run at 30 fps. For comparison, our implementation of Resnet-T, VQ-BeT, ACT and Diffusion Policy run at 100Hz, 100Hz, 50Hz, and 12Hz respectively.
@@ 5.6 Ablations
Vector Quantization: We replace the FSQ layer with a Vector Quantization layer of nearly the same codebook size , shown in the VQ column of Table 1. We see that FSQ’s superior codebook utilization leads to an improvement in performance.
@@
Observation Conditioned Decoder: Many prior condition the action decoder with current observation [76, 30]. We experiment with this by appending observation tokens to the skill tokens and allowing the transformer decoder to jointly cross-attend to both , shown in Table 1. We see that conditioning on observations leads to a deterioration in performance.
@@
Mirrored Decoder: Following a typical autoencoder design, we use a decoder that mirrors the encoder, using transposed convolutions instead of strided convolutions, and with the strides in reverse order as in the encoder. This decoder directly takes skill-token embeddings as input and outputs the continuous actions , and results are shown in Table 1. We see that this method performs worse, suggesting attending to all quantized codes in , as our decoder does, is important for faithfully predicting actions.
@@
Causality: We ablate the use of causal layers in various parts of our network in Table 2. We see that removing causality from any part of our architecture leads to worse performance, suggesting that causal masking imparts the model with a helpful inductive bias for modeling robot action data.
We also perform a sensitivity experiment over several hyperparameters including downsampling factor and codebook size in Figure 4. Across the board we see that the hyperparameters are more important in the difficult few-shot learning setting. In Figure 4(a) we see that both algorithms have the best performance with a modest downsampling factor of , and in Figure 4(b) we see that QueST does well with a 1024 codebook vectors. For more discussion on ablations please refer Appendix LABEL:app:abldiscussion
5.7 Latent Skill-Space Analysis
We present a t-SNE visualization (Figure 5) illustrating the learned skill-space across multiple set of similar tasks. We consider four different combinations of similar tasks to effectively examine the z-embeddings corresponding to their trajectories. Each data point in the plot represents a vector of z-embeddings at a specific timestep throughout the entire episode, with decreasing transparency indicating temporal progression. We show that the QueST encoder learns a semantically meaningful skill-space that encodes shared representations of similar motion primitives across different tasks. This analysis includes the first 11 tasks from LIBERO-90. For better comprehension, we encourage readers to review the corresponding rollouts on the website. Notably, the skill-space learning happens in the first stage training which does not make use of any task labels.
@@ 6 Conclusion
This paper works towards the broader goal of automating a wide range of manipulation tasks. While this can have positive impacts, such as helping people with mobility impairments or performing menial tasks humans would rather not do, it can also have negative impacts such as automating peoples’ jobs away and further concentrating wealth in the hands of a handful of companies. It is important that we in the machine learning community advocate for equitable use of the technology we develop.
@@ B.3 Baseline Implementation:
To ensure fair comparison of different model architectures, we use same input modalities and same observation & task encoders for all baselines , for both LIBERO and Metaworld experiments. VQ-BeT needs a goal image, we instead give it task embedding as goal. Same as QueST, we concatenate observation embeddings for all modalities at any timestep and project them to respective model’s hidden dimension.
@@ B.4 Compute:
B.5 Discussion on Baselines
Many recent works use discrete latent variable models as a mechanism to learn shared abstractions over continuous low-level skills. QueST, VQ-BeT and PRISE all do this and perform two- staged learning. However, in this work we propose several important architecture choices leading to QueST’s strong performance. Specifically, QueST encodes actions to a sequence of n encodings (skill tokens) using a novel autoencoder that captures temporal correlation within an action sequence with causal convolution and masked self-attention layers.
Like QueST, VQ-BeT performs temporal abstraction by encoding a sequence of actions into one state-independent latent vector. It concatenates input actions and encodes the sequence to just one single latent encoding using an MLP which does not explicitly model any temporal correlation. A sampled action sequence might contain multiple motion primitives of variable length and start point, and capturing them with a single encoding is limiting as it restricts abstraction at different levels of granularity. This is validated by the poorer () performance of VQ-BeT with a larger chunk size of . QueST flexibly captures this variability within encodings using its encoder, specifically designed to model temporal correlation within input actions. With an effective codebook size of , QueST’s effective latent space is while that of VQ-BeT is . Its relatively small latent space limits its expressive capacity, and thus limits its ability to learn sharable representations between tasks, as evidenced by its worse few-shot performance and reliance on continuous offset prediction (whereas QueST can achieve high success rate using only the output from its action decoder).
PRISE performs temporal abstraction by learning discrete codes for state-action pairs and using BPE to group common token sequences into higher-level skills. However, BPE is known to suffer with evolving language leading to a suboptimal character-level tokenization, and it might struggle to effectively encode new action sub-sequences that the model has not seen before (eg. stitching sequences in between two tasks in LIBERO-LONG). This is primarily why decoder finetuning is necessary in PRISE. On the contrary, QueST is more end-to-end as it lets the encoder handle temporal abstraction in the encoding phase itself and gracefully encodes new action-sequences as combinations of its learned latent codes. Our few-shot results without decoder-finetuning shows that such a unified approach learns more generalized abstractions than the baselines and can more effectively represent new action sequences in unseen tasks.
Despite tuning the hyperparameters, the multitask performance of ACT is very low (54%) on LIBERO. Moreover, as per BAKU [26], ACT’s performance drastically reduces when number of demos are reduced to 35-per-task in MetaWorld. This suggests ACT is very data-hungry and hence struggles in low-data regime. Another major issue with ACT is tuning the action chunk size to which the method is very sensitive (-14% in metaworld fewshot with chunk size of 32). While we report QueST results for chunk size 32, we did observe very little variation (<1% in LIBERO) with size 16,48 and 64, indicating robustness to this hyperparameter.
@@ Appendix C Discussion on Ablations
Appendix D Additional Results
D.1 Fewshot IL
@@
In finetuning phase, we finetune ResNet-T, VQ-BeT & QueST for 100 epochs and ACT & Diffusion Policy for 200 epochs. For each task in MetaWorld, we evaluate each method across 10 evenly spaced checkpoints for 5 seeds on 50 distinct initial states and report the results corresponding to the best performing checkpoint. For Libero, we found the final checkpoint to perform best for all methods and hence report results corresponding to it across 9 seeds.